OpenStudy (anonymous):

The value of n-3 is how much less than the value of n+3. A. 3 B. 6 C. n+3 D. n-6 E. n I don't get what the difference is between B and D?

OpenStudy (anonymous):

(n+3)-(n-3)=n+3-n+3=6

OpenStudy (anonymous):

answer is B

OpenStudy (anonymous):

Yes. But why isn't it D?

OpenStudy (anonymous):

Because you can substitute in a number for n to check. What happens when you do that?

OpenStudy (anonymous):

Oh it would 6 less than n instead of 3 less in answer B. Thanks for clearing that up.

OpenStudy (anonymous):

The terms n are cancelled out as i already proved the solution